OpenStudy (anonymous):

divide. write answer using nonnegative exponents 6x^4y^2/2x^7t^2

OpenStudy (saifoo.khan):

wht about the last one?

OpenStudy (anonymous):

I just moved forward. I can always go back. I just need to get it done.

OpenStudy (saifoo.khan):

Oh, okay!

OpenStudy (saifoo.khan):

its correct.

OpenStudy (anonymous):

ok what about this one?

OpenStudy (saifoo.khan):

i ca do this too

OpenStudy (anonymous):

can or can't?

OpenStudy (saifoo.khan):

can* lol.

OpenStudy (anonymous):

ok. please! :)

OpenStudy (saifoo.khan):

3y^2 ------- x^3 t^2

OpenStudy (anonymous):

that says incorrect

OpenStudy (anonymous):

see here: 6x^4y^2/2x^7t^2 here according to law of exponents a^b^c = a^bc so 6x^4y^2 = 6x^(4y*2) = 6x^8y and now in denominator we have 2x^7t^2 = 2x^(7t*2) = 2x^14t so we got (6x^8y)/(2x^14t) is it the answer

OpenStudy (anonymous):

may be i had done wrong something if it is also wrong but if u wait for 1 second i would answer u

OpenStudy (saifoo.khan):

kush, please check her last prob too.

OpenStudy (anonymous):

it says yours is wrong too.

OpenStudy (saifoo.khan):

@Rachel, im heading to sleep now. Dad's here. Bye! Have a nice time.

OpenStudy (anonymous):

ok thanks

OpenStudy (saifoo.khan):

please chk that is t?

OpenStudy (anonymous):

it is